Eliminations Debit Consolidated Paulee Credit Sergio INCOME STATEMENT Sales Invest. inoome from Sergio $ 43,000 $20,000 f $6, this is the answers i need help figuring out how they got the numbers thank you. !
12) Paulee Corporation paid $24,800 for an 80% interest in Sergio Corporation on January 1, 2013, at which time Sergio's stockholders' equity consisted of $1 5,000 of Common Stock and $6,000 of Retained Earnings. The fair values of Sergio Corporation's assets and liabilities were identical to recorded book values when Paulee acquired its 80% interest. Sergio Corporation reported net income of $4,000 and paid dividends of $2,000 during 2013. Paulee Corporation sold inventory items to Sergio during 2013 and 2014 as follows: 2013 Paulee's sales to Sergio Paulee's cost of sales to Sergio Unrealized profit at year-end 1,000 1,500 2014 $5,000 $6,000 3,000 3,500 At December 31, 2014, the accounts payable of Sergio include $1,500 owed to Paulee for inventory purchases. Required: Financial statements of Paulee and Sergio appear in the first two columns of the partially completed working papers. Complete
the consolidation working papers for Paulee Corporation and Subsidiary for the year ended December 31, 2014. Packo Paulee Corporation and Subsidiary Consolidation Working Papers for the year ended December 31, 2014 Eliminations Debit Paulee Sergio Credit Consolidated INCOME STATENENT Sales Invest. income from Sergio $ 43,000 $20,000 6,700 Cost of Sales other expenses Noncontrolling interest share Net income (22,000) (12,200) 8.000) 3,000) $ 15,500 $9,000 Retained Earnings 1/1 Add: Net income Less: Dividends Retained Earnings 12/31 BALANCE SHEET Cash Net Receivables Dividend Receivable $9,000 15, 500 (10,000) S 8,000 9,000 ( 5,000) $ 14,500 $12,000 $3,000 10,000 $ 5,400 14,000 2,000 18,000 8,000 Inventories Goodwill Plant assets-net 24,000 31,000 Investment in Sergio TOTAL ASSETS LIAB. & EQUITY Accounts payable 17, 500 Dividend payable Other debt Capital stock Retained Earnings 1/1 Noncntrl. 28,100 $ 91,500 $52,000 $12,500 2,500 10,000 15,000 7,000 12,500 40,000 12,000 14,500 Interest 12/31 Nonentrl. Interest TOTAL LIAB. EQUITY $52,000 $91,500
Eliminations Debit Consolidated Paulee Credit Sergio INCOME STATEMENT Sales Invest. inoome from Sergio $ 43,000 $20,000 f $6,000 $57,000 6,700 b 6,700 6,000 1,000 f Cost of Sales Other expenses Nonoontrolling interest share Net income (22,000) ( 12,200) 8,000) g 3,000) (24,500) (15,200) 1,500 a 1,800) 1,800 $ 15,500 $9,000 $15,500 Retained Earnings 1/1 Add: Net income Less: Diidends $ 9,000 15,500 (10,000) S 9,000 15,500 1,000 10,000) $ 8,000 9,000 5,000) 8,000 4,000 Retained Earnings 12/31 BALANCE SHEET Cash Net Reoeivables Dividend Receivable Inventories Goodyil1 $ 14,500 $14,500 $12,000 $ 8,400 22,500 $5,400 14,000 S3,000 10,000 1,500 2,000 2,000 18,000 1,500 24,500 10,000 55,000 8,000 a 10,000 24,000 31,000 Plant assets-net 2,700 26,400 1,000 b Investment in d Sergio 28,100 $120,400 $ 91,500 $52,000 TOTAL ASSETS LIAB. & EQUITY Accounts payable $17, 500 Dividend payable Other debt Capital stook Retained Earnings 1/1 Nonontrl. Interest 12/31 Nonontrl. $28,500 7,500 22.500 1,500 2,000 $12,500 2,500 10,000 15,000 7,000 12,500 40,000 40,000 15,000 14,500 12,000 14,500 6,600 800 7,400 $120,400 Interest $52,000 $ 91,500 TOTAL LIAB.& SQUITY Objective: LO5.3 Defer unrealized inventory profits remaining in the ending inventory Difficulty: Difficult AACSB: Application of knowledge b
Answer #1

Firstly the consildated statements is for the year ended Dec 2014.Therefore impact of all the adjustment for the mentioned period is taken into consideration.

Secondly , the basis on which the numbers are provided in the answer is describe hereunder:-

1.As mentioned in the question, in 2014 there is inter company sale of Rs 6000 by Paulee to Sergio-As consolidated Financial statements for both the companies are made therefore its is necessary to eleminate Rs 6000 to record the correct Revenue(Sale).

2. 6700 being income received by Paulee from Sergio for holding interest in the company. The amount of 6700 which is income for Paulee is an Appropriation for Sergio. Therefore it is necessary to eleminate the Figure to report correct Consolidated Statments for the Dec 2014.

3. a)Unrealised Profit of $ 1500 in the book of Paulee for sale of inventory should be eleminated ,as in relaity there is no such profit earned if consolidated statements are prepared.

b) $ 6000 and $1000 should be deducted form the cost of sales as already inter company sale of $ 6000 is elemintaed from the books as per point 1. above and the same should not form part of purchase and the unrealiesed profit carried forward which was initally taken into account should be eleminated from cost of sales as there is no such profilt earned in consolidated statements.

4.Paulee holds 80% controlling interest in Sergio.The remaming 20% is held by outsiders.$ 9000 is the net income of Sergio for Fy 2014.

20% of 9000 is equal to 1800 is share odf outsiders which is elmianted to arrive at the correct income earned by the companies.

5.$ 8000 of retained earinings has been eleminated as the impact is already been provided and in consolidated statements its necessary to eleminate.

6. $5000 total divided paid:-

20% of 5000=1000 is dividend paid for non controlling interest which will have no impact in conolidated statements therefore need to be eleminated.

and the remaning 80% of 5000=4000 is the dividend recived by Paulee which again should be eleminated as it is appropraiation for sergio and cross adjustments should be eleminated.

7.Net Receivable of Paulee includes $ 1500 from Sergio for sale on Inventroy.As this is the cross adjustment the amount due from sergo will be appreaing in sergios Accounts payable and the same is included in net receviable of Paulee.The amount is leminated as this is a cross adjustment.

8.$2000 dividend receivable by Paulee is the amount which is Payable By Sergio >therefor it is eleminated as both debit and credit sets off as consolidated Statments are prepared.

9.$1500 is the unrealised Profit which need to be eleminated from the amount of Inventory as no such profit is earned in relaity.and for preparing Consolidated Statments the amount needs to be eleminated otherwise inventory will be overpriced by $ 1500 whereas in relaity there is no such profit earned.

10.Paulee acquired 80% controlling interest for$ 24800. Therefor 100% interest is of =24800/80*100=$31000

Retained Earning + Stock =15000+6000=21000

Therefore Goodwill=31000-21000=$10000

11.$1000 is added as it is unrealised profit of the last year whose impact is in Financial statement of Paulee

12.2700 is net increment in the vcalue of investment in sergio shares which will be elemintaed

B=26800 is the total of 24800 paid for controlling interest + 1600 for the dividend recevied in 2013.

13.$ 1500 accounts Payable is a cross adjustment having same amount diue to pAulee as Receviable.Therfore there would be overstating of Payables and needs to be eleminated in preparing consolidated statements.

14.$ 2000 Dividend Payable by Sergio is receivable By paulee again it has no impact as its a cross adjustments and need to be eleminated to derive the correct figure of dividend Payable.

15.Captal Stock of $ 15000 is eleminated as it is already included in Pauless Books.

16.$6600 is calculated as follows:-

Total 100\% Share is of $ 31000. Therefore , 20% of 31000=6200

also add dividend received in 2013 for non controlling interest i.e 20% of 2000=400

Total amount of Liability =6200+400=6600

Calculation of $ 800 is total earning in 2014 is 9000 out of which 5000 is distributed as dividend .Therefore net income after distribution of dividend is =9000-5000=4000

20% is non controlling interest which should be added as an liabilty therefore 20% of 4000=800 is net liability .